Module 10 Profit Worksheet

1. Fill out the productivity chart for Spicy Chicken sandwiches at each level of labor at Chick-fil- B:

Workers Output Average Product Marginal Product

0 0

1 15

2 35

3 45

2. Now suppose that Chick-fil-B has to pay $100 per day to turn the lights on and $8 per hour for the workers. Complete the associated table and suppose that each worker just works one hour.

Workers Output Fixed Cost Variable Cost Total Cost Marginal Cost

0 0

1 15

2 35

3 45

3. Calculate the average fixed cost, average variable cost, and average total costs.

Workers Output   Average Fixed Cost Average Variable Cost Average Total Cost

0 0

1 15

2 35

3 45

4. Suppose that Chick-fil-B receives $5 for every spicy chicken sandwich that they sell. Complete the associated table.

Workers Output Total Revenue Marginal Revenue Total Cost Marginal Cost  

0 0

1 15

2 35

3 45

5. What is the highest profit?
Answer #1

Question1


marginal product Workers output O O 11 15 2 35 3 45 average product O 15 17.5 15 10

Average product = output / worker

For worker 1=15/1 =15.

Same calculation will be done for further.

Marginal product =change in quantity / change in number of workers

For worker 1 =15-0/1-0 =15/1 =15.

Same will be calculate further.

Question 2.

variable Total cost marginal cost cost Fixed Worker output cost 10 100 15 100 35 100 3 45 100 8 100 108 0.53 0.4 116 ||116 24

Fixed cost= if workers work or not but 100$ is given so it is called fixed cost. Fixed cost remains same and never change with increase of worker.

Variable cost = here $8 per worker cost is given and that is variable cost.

Variable cost = no of worker *per hour wage

For 1 worker =1*8

=8

Same will beccalculate further

Total cost = fixed cost+ variable cost   

For 1 worker= 100+8 =108

Same will be calculate further.

Marginal cost= change in cost ÷ change in quantity

For 1 worker = 108-100/15-0 =8/15 =0.53

Same will be calculate further.

Question 3


AFC C Worker o 1 |2 output o [15 35 AFC o |6.67 |2.85 |2.22 AVC o |0.53 |0.46 ATC || 7.2 |3.31 2.75 45 |0.53.

AFC= FIXED COST /QUANTITY

AVC= VARIABLE COST /QUANTITY

ATC=TOTAL COST /QUANTITY

Question 4.


total Marginal Worker output revenue revenue O 1 2 3 O 15 35 45 O 75 175 225 | total marginal cost cost 100 108 0.53 116 0.4

Revenue =price *quantity

For 1 worker=5*15 =75

Same calculation will be done further.

Marginal revenue = change in total revenue / change in quantity

For 1 worker= 75-0/15-0 =5.

Same calculation will be done further.

Total cost and marginal cost are brings from Question 2 answer.

Question 5.

Profit =total revenue - total cost

For worker 0= 0-100= - 100(loss)

Worker 1= 75-108 =-33(loss)

Worker 2= 175-116= 59(profit)

Worker 3= 225-124= 101(profit)

So highest profit is 101